OpenStudy (anonymous):

Solve the following system of equations x+9y=9 (1) x=5-9y (2)

hero (hero):

5-9y = 9 - 9y no solutions

OpenStudy (saifoo.khan):

x+9y=9 x=5-9y simplify.. x+9y = 9 x+9y = 5 since their slope is same, but the intercepts are different, so it has NO SOLUTIONS!

OpenStudy (anonymous):

\[5\neq 9\]

OpenStudy (anonymous):

thank you all

OpenStudy (saifoo.khan):

uw

OpenStudy (anonymous):

NO SOLUTION.